OpenStudy (anonymous):

which ordered pairs is not a solution of y=2x+1 (3,7) (0,1) (-1,1) (-3,-5)

OpenStudy (anonymous):

plug them in

OpenStudy (anonymous):

I don't know how. can you explain it to me.

OpenStudy (anonymous):

coordinate points are ( x , y ) correct? just substitute

OpenStudy (anonymous):

okay what I mean is explain the whole thing to me

OpenStudy (anonymous):

on the coordinate plane there is a x horizontal y vertical 

OpenStudy (anonymous):

how u name a coordinate point on the graph ( 2 , 2 ) coordinate point ^ ^ | | ( x , y ) y = m x + 0 ^ ^ 2 2 2 = 2 + 0 2 = 2 ( 2 , 2 ) it is a solution to the equation